OpenStudy (anonymous):

If f(x) = 4^-x, then f(a/2) = (a) 2 (b) 1 (c) 2^1/a (d) 1/2^a (e) 2^a

Parth (parthkohli):

Plug and play again!\[f\left({a \over 2}\right) = 4^{-{a \over 2}} \]

OpenStudy (anonymous):

does that equal to 1/4^a/2 ?

Parth (parthkohli):

Yes sir :)

OpenStudy (anonymous):

Could be wrong, but I think it's one of those... Closest is (c)... f(a/2)=4^(-a/2)=[4^(1/2)]^-a=2^-a

Parth (parthkohli):

You could simplify that.

OpenStudy (anonymous):

1/square root 4^a

Parth (parthkohli):

Yes, and?

OpenStudy (anonymous):

are u supposed to rationalize it?

Parth (parthkohli):

Yes sir!

OpenStudy (anonymous):

so what do u do after square root 4^a/ 4^a

Parth (parthkohli):

Remember that you have\[(4^a)^{1 \over 2} \over 4^a \]Subtract exponents.

Parth (parthkohli):

lol wait we are back at the point that we started with.

OpenStudy (anonymous):

4^1/2 is 2 but how did it change do 2^-a?

Parth (parthkohli):

Let's check.

Parth (parthkohli):

\[ 4^{-a \over 2} \implies (4^{-a})^{1 \over 2}\implies (2^{-2a})^{1 \over 2} \implies 2^{-a}\]

Parth (parthkohli):

That was quick.

OpenStudy (anonymous):

\[4^{-\frac{a}{2}}=\frac{1}{4^{\frac{a}{2}}}=\frac{1}{\sqrt{4}^a}=\frac{1}{2^a}=2^{-a}\] how'd i do?

OpenStudy (anonymous):

oh not so good because i didn't match it up

Parth (parthkohli):

You did well, son.

Parth (parthkohli):

;P

OpenStudy (anonymous):

\[\frac{1}{2^a}=\left(\frac{1}{2}\right)^a\]

OpenStudy (anonymous):

so is the answer (d) ?

Parth (parthkohli):

Yes. lawls.

OpenStudy (anonymous):

thanks (y)